There are 30 good computers and 6 defective ones. In how many ways can 18 computers be selected so that exactly 2 computers are defective?
Answer by Edwin McCravy(20063)

There are 30 good computers and 6 defective ones. In how many ways can 18 computers be selected so that exactly 2 computers are defective?
"30 choose 16" TIMES "6 choose 2"

       C(30,16)×C(6,2) = 145422675×15 = 2181340125 ways.

Edwin